A 1.764-g sample of heptanoic acid, C7H14O2 (130.19 g/mol) was burned in a bomb calorimeter with excess oxygen. The temperature of the calorimeter and the water before combustion was 23.68 °C; after combustion the calorimeter and the water had a temperature of 32.12 °C. The calorimeter had a heat capacity of 500 J/K, and contained 1.462 kg of water. Use these data to calculate the molar heat of combustion (in kJ) of heptanoic acid.
Answer #1

Following is the - complete Answer -&- Explanation: for the given: Question: in....typed format....

\RightarrowAnswer:

Molar heat of combustion of heptanoic acid ( C7H14O2 ) : Q =   4138.52 kJ /mol  ...

\RightarrowExplanation:

Following is the complete Explanation: for the above Answer...

  • Given:
  1. ​​​​​​​Mass of heptanoic acid : ( C7H14O2 ) : m1 = 1.764 g ( grams )
  2. Molar mass of heptanoic acid : ( C7H14O2 ) : n1 = 130.19 g/mol ( grams per mole )
  3. Temperature of the calorimeter & water: before combustion (initially ): Ti =   23.68 oC =  ( 273.15 + 23.68 ) = 296.83 K ( Kelvin )
  4. Final temperature of the calorimeter and water: Tf = 32.12 oC = ( 273.15 + 32.12 ) = 305.27 K ( kelvin)
  5. Mass of water in the calorimeter: mh2o = 1.462 kg ( kilo-grams )
  6. Heat capacity of the calorimeter: Ccal = 500.0 J / K = 0.500 kJ / K ( kilo-joule per kelvin )
  7. We know: Specific heat of Water: Ch2o = 4.186 kJ / kg. K
  • ​​​​​​​Step - 1:

\Rightarrow Number of moles of  heptanoic acid : ( C7H14O2 ) : =  m1 / n1 =  ( 1.764 g ) / ( 130.19 g/mol ) = 0.0135 mol

  • Step - 2:

​​​​​​​\Rightarrow Change in temperature of the calorimeter and water : \DeltaT =  Tf - Ti = ( 305.27 - 296.83 ) = 8.44 K ( Kelvin )

\Rightarrow Heat gained by calorimeter: Qcal = Ccal x  \DeltaT =  ( 0.500 kJ/K ) x ( 8.44 K ) = 4.22 kJ

\Rightarrow Heat gained by water, in caloriemter: Qh2o = mh2o x Ch2o x  \DeltaT

= ( 1.462 kg ) x (4.186 kJ / kg. K ) x ( 8.44 K )

= 51.65 kJ ( kilo-joule )

  • Step - 3:

Therefore: total heat of combustion: ( Qcomb ) will be equal to the following:

\Rightarrow    Qcomb =  Qcal + Qh2o =  ( 4.22 + 51.65 ) kJ = 55.87 kJ   ( kilo-joule )

  • Step - 4:

​​​​​​​As we can see above, the following : i.e.

\Rightarrow 0.0135 mol ( moles ) of heptanoic acid : ( C7H14O2 ) , after combustion gives: 55.87 kJ ( kilo-joule ) of Heat ...

\Rightarrow1.0 mol ( moles ) of heptanoic acid : ( C7H14O2 ), would give: ( 55.87 kJ ) / ( 0.0135 mol ) = 4138.52 kJ ( approx. ) of heat ...

Therefore:

\RightarrowMolar heat of combustion of heptanoic acid : ( C7H14O2 ) =  4138.52 kJ /mol